Your daughter is born today and you want her to be a millionaire by the time she is
35 years old. You open an investment account that promises to pay 12% per year.
How much money must you deposit each year, starting on her 1st birthday and
ending on her 35th birthday, so your daughter will have RM1,000,000 by her 35th
birthday?
1
Expert's answer
2015-03-12T10:30:38-0400
n = 35 years, r = 12% per year.
To calculate the annual deposit amount, we use the formula of compound interest:
FV = PV*(1+r)^n
PV = 1000000/(1.12 + 1.12^2 + … + 1.12^35) = $2,068.41 per year.